# Resolution of Simpson’s paradox via the common cause principle

### v1

@@ Abstract
Simpson’s paradox is an obstacle to establishing a probabilistic association between two events $a_{1}$ and $a_{2}$, given the third (lurking) random variable $B$. We focus on scenarios when the random variables $A$ (which combines $a_{1}$, $a_{2}$, and their complements) and $B$ have a common cause $C$ that need not be observed. Alternatively, we can assume that $C$ screens out $A$ from $B$. For such cases, the correct association between $a_{1}$ and $a_{2}$ is to be defined via conditioning over $C$. This set-up generalizes the original Simpson’s paradox. Now its two contradicting options simply refer to two particular and different causes $C$. We show that if $B$ and $C$ are binary and $A$ is quaternary (the minimal and the most widespread situation for valid Simpson’s paradox), the conditioning over any binary common cause $C$ establishes the same direction of the association between $a_{1}$ and $a_{2}$ as the conditioning over $B$ in the original formulation of the paradox. Thus, for the minimal common cause, one should choose the option of Simpson’s paradox that assumes conditioning over $B$ and not its marginalization. For tertiary (unobserved) common causes $C$ all three options of Simpson’s paradox become possible (i.e. marginalized, conditional, and none of them), and one needs prior information on $C$ to choose the right option.
I Introduction
@@
Simpson’s paradox was discovered more than a century ago pearson ; yule , generated a vast literature, and is well-recognized in several fields cohen ; simpson ; blyth ; lindley ; barigelli ; barigelli2 ; pearl ; pearl2 ; hernan ; appleton ; armi ; agresti ; cohen_uncertainty . This counter-intuitive effect is said to limit one’s ability to draw obvious conclusions from probabilistic data. The effect is important because it demands more than simply extracting relative frequencies from data; e.g. it necessitates looking at exchangeability or causality lindley .
@@
where $B$ also occured in an earlier time: $t_{B}\leq t_{A_{1}}$. For example, $a_{1}$ and $\bar{a}_{1}$ can be positive and negative results, while $a_{2}$ and $\bar{a}_{2}$ mean two different medical treatments; $b$ and $\bar{b}$ can then refer to different features of patients, e.g. their sex ($b={\rm male}$, $\bar{b}={\rm female}$) or age ($b={\rm old}$, $\bar{b}={\rm young}$) lindley ; see Appendix A for more examples. Now since (2, 3) hold for each value of $B$ we should perhaps conclude that ${a}_{2}$ enables $a_{1}$ in contrast to (1). Decision-makers would not know whether to apply (1) or (2, 3). This is Simpson’s paradox. For the paradox, $A_{2}$ and $B$ must be dependent variables  [^1] We leave aside the following pertinent problem; see cohen_uncertainty for details. If probabilities are extracted from finite populations, the more conditioned version (2, 3) is less reliable, because it is extracted from a smaller population. For us all probability-providing populations will be sufficiently large. .
@@
Over time, several resolutions to the paradox have been proposed. Barigelli and Scozzafava barigelli ; barigelli2 proposed to replace (1) by $p(a_{2})p(a_{1}|a_{2})<p(\bar{a}_{2})p({a}_{1}|\bar{a}_{2})$, i.e. to interchanging $A_{1}$ and $A_{2}$ in (1). Then it is easy to see that its inversion under additional conditioning over $B$ is impossible. While (1) stands for prediction, i.e. aiming at $a_{2}$ (and not at $\bar{a}_{2}$) will more likely produce $a_{1}$ (than $\bar{a}_{1}$), the proposal by Ref. barigelli ; barigelli2 looks for retrodiction. Though retrodicting (in contrast to predicting) does not suffer from Simpson’s paradox, retrodicting and predicting are different things, hold different intuitions, and cannot generally be substituted for each other. In a related proposal, Rudas also sought to change the criterion (1) so that it does not allow inversion after additional conditioning over $B$, but still has several reasonable features rudas . The proposal is to employ $p(a_{2})[p(a_{1}|a_{2})-p(\bar{a}_{1}|a_{2})]<p(\bar{a}_{2})[p(a_{1}|\bar{a}_{2})-p(\bar{a}_{1}|\bar{a}_{2})]$ instead of (1) rudas . Notice the conceptual relation of this with the previous proposal  [^2] An unnatural point of both these proposals is that they depend on the ratio $p(a_{2})/p(\bar{a}_{2})$; e.g. for the example mentioned after (3) this means that the treatment that was applied more has better chances to be accepted. This drawback is also acknowledged in rudas . .
According to Lindley and Novick, the paradox may be resolved by going beyond probabilistic considerations (as we do here as well) and by employing the notion of exchangeability or causality lindley . Within that proposal, the data generally provides only propensities, and one needs additional assumptions of sample homogeneity (exchangeability) for equating propensities with probabilities even for a large sample size. They studied two examples [see Examples 1 and 2 in Appendix A], where natural assumptions on exchangeability lead to preferring (1) for Example 2 and (2, 3) for Example 1 lindley . Lindley and Novick also proposed that the same results can be found via using causality instead of exchangeability lindley . Pearl elaborated this proposal showing that the examples of Ref. lindley can be represented via time-ordered direct acyclic graphs (TODAG) pearl ; pearl2 ; see Fig. 1 and Appendix A for details. The suggestion is to condition over $B$ [hence using (2, 3)] if $B$ causes both $A_{1}$ and $A_{2}$. This is because conditioning over the cause reduces spurious correlations. In contrast, it is advised to use (1) if $B$ is an effect of $A_{2}$, but still a cause of $A_{1}$ pearl ; pearl2  [^3] The intuition of this suggestion is seen in the extreme case when $B$ screens $A_{1}$ and $A_{2}$ from each other, i.e. $A_{1}$, $B$ and $A_{2}$ form a Markov chain. Then the conditional probability $p(A_{1}|A_{2},B)=p(A_{1}|B)$ will not depend on $A_{2}$ begging the original question in (1). This reasoning was generalized as the back-door criterion pearl . . For the two examples considered in lindley , Refs. pearl ; pearl2 make similar recommendations. The basis of these recommendations was criticized in armi .
Note, however, that the examples presented in Ref. lindley (and reproduced in Appendix A) do not need to support any TODAG. The data of these examples support an alternative interpretation that $B$ and $A=(A_{1},A_{2})$ do not have any causal influence on each other. There are also more examples of this type; see Appendix A. This is the starting point of our analysis: once $B$ and $A=(A_{1},A_{2})$ do not influence each other casually, it is natural to assume—following the common cause principle reich ; suppes —that there is an unobserved random variable $C$, which is a common cause for $A$ and $B$; see section II. An alternative assumption that leads to the same results is that $C$ causes only $A$, but it also screens out $A$ from $B$.
The introduction of this principle allows us to reformulate and extend Simpson’s paradox: (1) and (2, 3) —along with many other options—are particular cases that refer to particular choices of the unobserved cause $C$; see the next section. Now the paradox seems to be further from resolution than ever before. However, when $A_{1}$, $A_{2}$, $B$ and $C$ are binary (the minimal set-up of the paradox), and provided that the paradox holds (i.e. (1) and (2, 3) hold), all causes $C$ imply that the decision making is to be taken according to (2, 3); see the next section. The last section summarizes this work and discusses tertiary causes.

### v2

@@ Abstract
Simpson’s paradox is an obstacle to establishing a probabilistic association between two events $a_{1}$ and $a_{2}$, given the third (lurking) random variable $B$. We focus on scenarios when the random variables $A$ (which combines $a_{1}$, $a_{2}$, and their complements) and $B$ have a common cause $C$ that need not be observed. Alternatively, we can assume that $C$ screens out $A$ from $B$. For such cases, the correct association between $a_{1}$ and $a_{2}$ is to be defined via conditioning over $C$. This setup generalizes the original Simpson’s paradox: now its two contradicting options refer to two particular and different causes $C$. We show that if $B$ and $C$ are binary and $A$ is quaternary (the minimal and the most widespread situation for the Simpson’s paradox), the conditioning over any binary common cause $C$ establishes the same direction of association between $a_{1}$ and $a_{2}$ as the conditioning over $B$ in the original formulation of the paradox. Thus, for the minimal common cause, one should choose the option of Simpson’s paradox that assumes conditioning over $B$ and not its marginalization. The same conclusion is reached when Simpson’s paradox is formulated via 3 continuous Gaussian variables: within the minimal formulation of the paradox (3 scalar continuous variables $A_{1}$, $A_{2}$, and $B$), one should choose the option with the conditioning over $B$.
{IEEEImpStatement}
Simpson’s paradox is a genuine decision-making problem, and not just an artefact of a deficient data-gathering process. This also follows from our re-estimation of its frequency in an unbiased data generation process. Hence, the paradox can lead to incorrect recommendations in various fields ranging from artificial intelligence to psychology. This paper shows that the paradox can be resolved if the minimal (not necessarily observed) common cause (or screening variable) can be assumed for the involved random variables. The result holds both in discrete and continuous settings, which relates to linear regressions. On several non-trivial examples from the published literature (smoking, COVID-19, etc), we show that such an assumption is plausible and leads to new predictions. Given deep relations between the common cause principle and basic unsupervised learning methods (Non-negative Matrix factorization, Probabilistic Latent Dirichlet indexing), we expect that our results will facilitate connections between decision-making and machine learning algorithms.
{IEEEkeywords}
causality, confounding, decision making, probabilistic association
1 Introduction
[…]
@@ 1 Introduction
Simpson’s paradox was discovered more than a century ago [1, 2], generated a vast literature, and is well-recognized in several fields including, statistics, epidemiology, psychology, social science, etc. [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. This counter-intuitive effect limits the ability to draw conclusions from probabilistic data. The effect is important because it demands more than simply extracting relative frequencies from data; e.g. it necessitates looking at exchangeability [9] or causality [7, 8, 9, 13, 14].
The paradox starts with two random variables $A$ and $B$. Now $A=(A_{1},A_{2})$ contains control variable $A_{2}$ and the target variable $A_{1}$, while $B$ is a side random variable that correlates with both $A_{1}$ and $A_{2}$. The meaning of $A$ and $B$ is clarified via examples presented below. If there is no information on the outcome of $B$, the behavior of $A$ can be studied on two levels. The first (aggregated) level is that of marginal probabilities $p(A=a)$. The second level is finer-grained and is represented by conditional probabilities $p(A=a|B=b)$ for all possible values of $B$. Simpson’s paradox amounts to certain relations between those probabilities; see section 2 for details. It states that no decision-making is possible, because conclusions drawn from probabilities on different levels contradict each other. Without Simpson’s paradox, decision-making can proceed at the aggregate level, because looking at the fine-grained level is either redundant or inconclusive. Thus, Simpson’s paradox first and foremost involves decision-making. Moreover, it demonstrates limitations of the sure-thing principle [5], a pillar of traditional decision making [25, 26, 27]. A recent review of the sure-thing principle (and its limitations other than Simpson’s paradox) can be found in Ref. [28]. Limitations of probabilistic decision-making are important for the modern artificial intelligence (probability models, uncertainty estimation, etc).
In section 2, Simpson’s paradox is defined in detail, and previous efforts to resolve it in several specific situations are reviewed and criticized. In particular, we show that while certain previous solutions of the paradox assumed the existence of (causally-sufficient) time-ordered directed acyclic graphs (TODAGs) that describe the 3 variables involved in the paradox, several important examples of the paradox need not support this assumption; see sections 2.2.3, 4 and 5. Based on the previous literature, we also argue in section 2 that Simpson’s paradox is sufficiently frequent when the probabilities of the involved variables are generated from the unbiased (non-informative) distribution, modeled via Dirichlet density. Hence this is a genuine decision-making paradox and not an artifact due to inappropriate data gathering.
Our proposal here is to look for the resolution of the paradox by assuming that—given two correlated variables $A$ and $B$—there is a random variable $C$ that makes $A$ and $B$ conditionally independent; i.e., $C$ screens out $A$ from $B$. Examples of Simpson’s paradox show that such a $C$ is frequently plausible, though it is normally not observed directly. In particular, $C$ is conceivable if the correlations between $A$ and $B$ are not caused by a direct causal influence of $B$ on $A$. Then the existence of $C$ is postulated by the common cause principle. (If correlations are caused by a causal influence of $A$ on $B$, Simpson’s paradox can formally exist, but factually it is absent because the decision is obviously to be taken according to the aggregated level.)
Introducing the screening variable $C$ allows us to reformulate and extend Simpson’s paradox: its two options—along with many other options—refer to particular choices of $C$; see section 3. Now the paradox seems to be further from resolution than before. However, we show that when the variables $A_{1}$, $A_{2}$, $B$, and $C$ holding the paradox are binary (the minimal set-up of the paradox), the decision-making is to be made according to the fine-grained probabilities, i.e., the paradox is resolved. Such a definite relation is impossible for a tertiary (or larger) $C$: now depending on $C$ all options of Simpson’s paradox are possible, e.g. the precise control of $C$ can be necessary for decision-making.
Next, we turn to Simpson’s paradox for continuous variables, which was discussed earlier than the discrete formulation [1]. It holds the main message of the discrete formulation. In addition, it includes the concept of the conditional correlation coefficient (only for Gaussian variables is the random-variable dependence fully explained by the correlation coefficient). The continuous formulation is important because it applies to big data [23, 24, 29], and because (statistically) it is more frequent than the discrete version [30]. The advantage of continuous Gaussian formulation is that the general description of the paradox under the common cause is feasible; see section 6. For this situation, we show conceptually the same result as for the discrete version: in the minimal (and most widespread) version of the paradox, the very existence of an (unobservable) common cause leads to preferring the fine-grained option of the paradox.
The rest of this paper is organized as follows. Section 2 is a short but sufficiently inclusive review of Simpson’s paradox and its resolutions proposed in the literature  [^1] Among the issues not addressed in this paper is the explanation of Simpson’s paradox using counterfactual random variables. This subject is reviewed in [6]. . It also discusses two basic examples for illustrating different aspects of the paradox; see section 2.2.3. In section 2.3, we review results about how frequent the paradox is and re-estimate its frequency within an unbiased data generation. In section 3 we reformulate Simpson’s paradox by assuming that there is a common cause (or screening variable) $C$ behind the three variables. Now $C$ need not be observable, since we show that it will be sufficient to assume that it exists and (provided that all variables are binary) Simpson’s paradox is resolved by choosing its fine-grained option. A similar conclusion is reached for Gaussian variables; see section 6. Section 4 considers published data from Ref. [16] on a case of smoking and surviving. This example is not easily treated via the existing methods. Still, we show that the existence of a common cause for this situation is plausible and that Simpson’s paradox can be studied via our method and leads to a reasonable result. Section 5 treats data on COVID-19, which was suggested in Ref. [31]. We demonstrate that an assumption of a plausible common cause points to different conclusions than in Ref. [31]. The last section summarizes and outlines future research directions.
2 Formulation of Simpson’s paradox and previous works
2.1 Formulation of the paradox for binary variables and its necessary conditions
@@
where $B$ also occured in an earlier time: $t_{B}\leq t_{A_{1}}$. Examples supporting (13) are studied below (sections 2.2.3, 4 and 5) and also in Appendix .1. Since (2, 3) hold for each value of $B$ we should perhaps conclude that ${a}_{2}$ enables $a_{1}$ in contrast to (1). Decision-makers would not know whether to apply (1) or (2, 3). This is Simpson’s paradox. Its equivalent formulation is when all inequalities in (13) are inverted  [^2] We leave aside the following pertinent problem; see [19] for details. If probabilities are extracted from finite populations, the more conditioned version (2, 3) is less reliable, because it is extracted from a smaller population. For us all probability-providing populations will be sufficiently large. .
For Simpson’s paradox (13) to hold, it is necessary to have one of the following two conditions:
$p(a_{1}|\bar{a}_{2},b)<p(a_{1}|{a}_{2},b)<p(a_{1}|\bar{a}_{2},\bar{b})<p(a_{1}|{a}_{2},\bar{b}),$ (4)
$p(a_{1}|\bar{a}_{2},\bar{b})<p(a_{1}|{a}_{2},\bar{b})<p(a_{1}|\bar{a}_{2},b)<p(a_{1}|{a}_{2},b).$ (5)
To find these relations, expand $p(a_{1}|a_{2})$ and $p(a_{1}|\bar{a}_{2})$ over the probabilities in (4, 5) [cf. (23, 24)], and note that e.g. $p(a_{1}|a_{2})$ is a weighted mean of $p(a_{1}|{a}_{2},b)$ and $p(a_{1}|{a}_{2},\bar{b})$. Given that (4) or (5) hold, Simpson’s paradox can be generated via suitable choices of $p(b|a_{2})$ and $p(b|\bar{a}_{2})$. For such choices, it is necessary that
$\left(p(b|a_{2})-\frac{1}{2}\right)\left(p(b|\bar{a}_{2})-\frac{1}{2}\right)<0,$ (6)
i.e., $A_{2}$ and $B$ must be dependent variables.
2.2 Attempts to resolve the paradox
2.2.1 Replacing prediction with retrodiction
Over time, several resolutions to the paradox have been proposed. Barigelli and Scozzafava [10, 11] proposed to replace (1) by
$p(a_{2})p(a_{1}|a_{2})<p(\bar{a}_{2})p({a}_{1}|\bar{a}_{2}),$ (7)
@@ 2.2.1 Replacing prediction with retrodiction
i.e. to interchanging $A_{1}$ and $A_{2}$ in (1). Then it is easy to see that its inversion under additional conditioning over $B$ is impossible. While (1) stands for prediction, i.e. aiming at $a_{2}$ (and not at $\bar{a}_{2}$) will more likely produce $\bar{a}_{1}$ (than ${a}_{1}$), the proposal by Ref. [10, 11] looks for retrodiction. Though retrodicting (in contrast to predicting) does not suffer from Simpson’s paradox, retrodicting and predicting are different things, hold different intuitions, and cannot generally be substituted for each other.
Rudas also sought to change the criterion (1) so that it does not allow inversion after additional conditioning over $B$, but still has several reasonable features [32]. The proposal is to employ $p(a_{2})[p(a_{1}|a_{2})-p(\bar{a}_{1}|a_{2})]<p(\bar{a}_{2})[p(a_{1}|\bar{a}_{2})-p(\bar{a}_{1}|\bar{a}_{2})]$ instead of (1) [32]. Notice the conceptual relation of this with the previous proposal (7).
An unnatural point of both these proposals is that they depend on the ratio $p(a_{2})/p(\bar{a}_{2})$; e.g. for the Example 1 mentioned below this means that if the treatment was applied more, it has better chances to be accepted. This drawback is acknowledged in [32].
2.2.2 Exchangeability and causality
According to Lindley and Novick, the paradox may be resolved by going beyond probabilistic considerations (as we do below as well) and by employing the notion of exchangeability or causality [9]. Within that proposal, the data generally provides only propensities, and one needs additional assumptions of sample homogeneity (exchangeability) for equating propensities with probabilities even for a large sample size. Exchangeability and the closely related notion of ergodicity remain influential in the current analysis of statistical problems exemplified by Simpson’s paradox [33]. Lindley and Novick studied the following two examples that support Simpson’s paradox (more examples are discussed in sections 4, 5, and in Appendix .1).
Example 1. Medical treatment [9]. $A_{1}=\{a_{1},\bar{a}_{1}\}$ (the target variable) is the recovery rate of medical patients: $a_{1}={\rm recovery}$, $\bar{a}_{1}={\rm no~{}recovery}$. $A_{2}=\{a_{2},\bar{a}_{2}\}$ refers to a specific medical treatment: $a_{2}={\rm treatment}$, $\bar{a}_{2}={\rm no~{}treatment}$. $B=\{b,\bar{b}\}$ is the sex of patients: $b={\rm male}$, $\bar{b}={\rm female}$. The times to which the random variables $A_{1}$, $A_{2}$ and $B$ refer clearly hold $t_{B}<t_{A_{2}}<t_{A_{1}}$.
Example 2. Plant yield [9]. $A_{1}=\{a_{1},\bar{a}_{1}\}$ (the target variable) is the yield of a single plant: $a_{1}={\rm high}$, $\bar{a}_{1}={\rm low}$. $A_{2}=\{a_{2},\bar{a}_{2}\}$ refers to the variety (color) of the plant: $a_{2}={\rm dark}$, $\bar{a}_{2}={\rm light}$. $B=\{b,\bar{b}\}$ refers to the height of the plant: $b_{1}={\rm tall}$, $\bar{b}={\rm low}$. The times hold $t_{A_{2}}<t_{B}<t_{A_{1}}$.
Lindley and Novick proposed that assumptions on exchangeability lead to preferring (1) for Example 2 and (2, 3) for Example 1 [9]. They also proposed that the same results can be found by using causality instead of exchangeability [9]. The same proposal was made earlier by Cartwright in the context of abstract causality [7, 8]. Pearl elaborated this proposal assuming that the above examples can be represented via time-ordered direct acyclic graphs (TODAG) [13, 14], where an arrow $\to$ represents the influence of an earlier variable to the later one; see Fig. 1 for details. If we follow this assumption, then—given the time constraints for the examples—each of them can be related to a unique TODAG:
${\bf Example~{}1}:~{}~{}~{}B\to A_{2}\to A_{1}\leftarrow B,$ (8)
${\bf Example~{}2}:~{}~{}~{}A_{2}\to B\to A_{1}\leftarrow A_{2}.$ (9)
In (8) the suggestion is to condition over $B$ [hence using (2, 3)] if $B$ influences both $A_{1}$ and $A_{2}$ [9, 13, 14]. This is because conditioning over the cause reduces spurious correlations. This reasoning was generalized as the back-door criterion [13].
In contrast, it is advised to use (1) in (9) since $B$ is an effect of $A_{2}$, but still a cause of $A_{1}$ [9, 13, 14]. The intuition of this suggestion is seen in the extreme case when $B$ screens $A_{1}$ and $A_{2}$ from each other, i.e. $A_{1}$, $B$ and $A_{2}$ form a Markov chain. Then the conditional probability $p(A_{1}|A_{2},B)=p(A_{1}|B)$ will not depend on $A_{2}$ begging the original question in (1). Thus, for the two examples considered in [9], Refs. [13, 14] make similar recommendations. The basis of these recommendations was criticized in [17].
2.2.3 Criticism
Realistically, Example 1 need not to support any TODAG. In fact, both arrows $B\to A_{2}$ and $B\to A_{1}$ are generally questionable: sex need not influence the selection of the treatment, $B\not\to A_{2}$ (unless the data was collected in that specific way), and many treatments are sex-indifferent, i.e. $B\not\to A_{1}$. For Example 1 it is more natural to assume that $B$ does not causally influence $A$. In such a situation, the common cause principle proposes that there is an unobserved random variable $C$, which is a common cause for $A$ and $B$ [34, 35]; see section 3.
Similar reservations apply to Example 2: now $A_{2}\to B$ is perhaps argued on the basis of color ($A_{2}$) being more directly related to the genotype of the plant, while the height ($B$) is a phenotypical feature. First, color-genotype and height-phenotype relations need not hold for all plants. Second (and more importantly), it is more natural to assume that the plant genotype influences both its color and height than that the color influences height. Hence the genotype can be a common cause for $A$ and $B$.
2.3 How frequent is Simpson’s paradox: an estimate based on the non-informative Dirichlet density
To estimate the frequency of Simpson’s paradox under fair data-gathering, we can try to generate the probabilities in (13) randomly in an unbiased way, and calculate the frequency of holding the paradox [36, 30]. The best and widely accepted candidate for an unbiased density of probabilities is the Dirichlet density, which is widely employed in statistics and machine learning [37, 38]; see Ref.[39] for a recent review. The Dirichlet probability density for $n$ probabilities $(q_{1},...,q_{n})$ reads:
${\cal D}(q_{1},...,q_{n}|\alpha_{1},...,\alpha_{n})=$
$\frac{\Gamma[\sum_{k=1}^{n}\alpha_{k}]}{\prod_{k=1}^{n}\Gamma[\alpha_{k}]}\prod_{k=1}^{n}q_{k}^{\alpha_{k}-1}\,\delta(\sum_{k=1}^{n}q_{k}-1),$ (10)
$\int_{0}^{\infty}\prod_{k=1}^{n}{\rm d}q_{k}{\cal D}(q_{1},...,q_{n}|\alpha_{1},...,\alpha_{n})=1,$ (11)
where $\alpha_{k}>0$ are the parameters of the Dirichlet density, $\delta(x)$ is the delta-function, and $\Gamma[x]=\int_{0}^{\infty}{\rm d}q\,q^{x-1}e^{-q}$ is the Euler’s $\Gamma$-function. Since ${\cal D}(q_{1},...,q_{n})$ is non-zero only for $q_{k}\geq 0$ and $\sum_{k=1}^{n}q_{k}=1$, the continuous variables themselves have the meaning of probabilities.
Many standard prior densities for probabilities are contained in (2.3); e.g., homogeneous ($\alpha_{1}=...\alpha_{n}=1$), Haldane’s ($\alpha_{1}=...\alpha_{n}=\alpha\approx 0$), Jeffreys ($\alpha_{1}=...\alpha_{n}=1/2$). For estimating the frequency of Simpson’s paradox, Ref. [36] employed homogeneous and Jeffreys prior.
For modeling a non-informative Dirichlet density we find it natural to take
$\alpha_{1}=...\alpha_{n}={1}/{n}.$ (12)
The homogeneity feature, $\alpha_{1}=...\alpha_{n}$ in (12) is natural for an unbiased density. The factor $\frac{1}{n}$ in (12) makes an intuitive sense, since $\alpha_{1}=...\alpha_{n}$ become homogeneous (non-informative) probabilities. Eq. (12) arises when we assume that the distribution of random probabilities is independent of whether they were generated directly from (2.3) with $n$ components, or alternatively from (2.3) with $nm$ components $\alpha_{1}=....=\alpha_{nm}$, and then marginalized. This requirement indeed leads to (12), as can be checked with the following feature of (2.3):
$\int_{0}^{\infty}{\rm d}q_{n-1}^{\prime}\,{\rm d}q_{n}^{\prime}\,\delta(q_{n-1}-q_{n-1}^{\prime}-q_{n}^{\prime})\,\times$
${\cal D}(q_{1},...,q_{n-2},q_{n-1}^{\prime},q_{n}^{\prime}|\alpha_{1},...,\alpha_{n})$
$={\cal D}(q_{1},...,q_{n-2},q_{n-1}|\alpha_{1},...,\alpha_{n-2},\alpha_{n-1}+\alpha_{n}).$ (13)
The message of (13) is that aggregating over two probabilities leads to the same Dirichlet density with the sum of the corresponding weights $\alpha_{n-1}$ and $\alpha_{n}$.
We estimated the frequency of Simpson’s paradox assuming that 8 probabilities $p(A_{1},A_{2},B)$ in (13) are generated from (2.3, 12) with $n=8$ (binary situation). This amounts to checking two relations (they amount to (13) and its reversal)
$[p(a_{1}|a_{2})-p(a_{1}|\bar{a}_{2})][p(a_{1}|a_{2},b)-p(a_{1}|\bar{a}_{2},b)]<0,$ (14)
$[p(a_{1}|a_{2},b)-p(a_{1}|\bar{a}_{2},b)][p(a_{1}|a_{2},\bar{b})-p(a_{1}|\bar{a}_{2},\bar{b})]>0.$
Our numerical result is that the frequency of two inequalities in (14) is $\approx 4.29\%\pm 0.001\%$. For this precision it was sufficient to generate $N=10^{7}$ samples from (2.3, 12) with $n=8$. This result compares favorably with $\approx 1.66\%$ obtained for $\alpha_{1}=...\alpha_{8}=1$ (homogeneous prior), and $\approx 2.67\%$ obtained for $\alpha_{1}=...\alpha_{8}=0.5$ (Jeffreys prior) [36]. It is seen that the frequency of Simpson’s paradox is a decreasing function of $\alpha_{1}=...=\alpha_{8}=\alpha$ [36].
Roughly, the above result $\approx 4.29\%$ means that in every 1000 instances of 3 binary variables, 42 instances will show Simpson’s paradox. This number is reassuring: it is not very large meaning that the standard decision-making based on the marginal probabilities in (1) will frequently be reasonable. But it is also not very small, showing that Simpson’s paradox is generic and has its range of applicability.
3 Common cause principle and reformulation of Simpson’s paradox